25. Find the first five terms of the recursive sequence an = n + 3an-1 where a = -2.
Answer #1

25 an = ntzant a = -2, given. az=.2+ 30.2-1 . = 2 + 324 = 2+ 3(-2) = 2-6 = -4. ag = 3 + 3a3-1 = = 3+ 3 a 3.+ 3.(-4) = 3-12 =